# Key Words.: cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD1 Introduction

### v1

@@ Acknowledgements.
The first two authors should be regarded as joint first authors. We are grateful to S. Trager and A. Vazdekis for sharing with us their quality flags of the XSL and MILES libraries, respectively. We thank A.S. Bolton and M.Bernardi for useful discussions regarding SDSS velocity dispersions. We thank E. Buckley-Geer, M. Millon, and all the friends of the TDCOSMO collaboration for useful feedback that improved this manuscript. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1974. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. Based in part on observations collected at the European Southern Observatory data obtained from the ESO Science Archive Facility. Some of the data presented here were taken from the Sloan Digital Sky Survey Archive. Support for this work was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51492 awarded to AJS by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. We acknowledge support by NSF through grants NSF-AST-1906976 and NSF-AST-1836016, and from the Moore Foundation through grant 8548. We acknowledge support from NASA through grants JWST-GO-1974 and JWST-GO-2974. SB thanks the Department of Physics and Astronomy, Stony Brook, for their support.References
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35
- Bacon et al. (2010) Bacon, R., Accardo, M., Adjali, L., et al. 2010, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 7735, Ground-based and Airborne Instrumentation for Astronomy III, ed. I. S. McLean, S. K. Ramsay, & H. Takami, 773508
- Bacon et al. (2010) Bacon, R., Accardo, M., Adjali, L., et al. 2010, in SPIE Conference Series, Vol. 7735, Ground-based and Airborne Instrumentation for Astronomy III, ed. I. S. McLean, S. K. Ramsay, & H. Takami, 8
- Barth et al. (2002) Barth, A. J., Ho, L. C., & Sargent, W. L. W. 2002, The Astronomical Journal, 124, 2607
- Birrer et al. (2024) Birrer, S., Millon, M., Sluse, D., et al. 2024, Space Sci. Rev., 220, 48
- Birrer & Treu (2021) Birrer, S. & Treu, T. 2021, A&A, 649, A61
- Birrer et al. (2019) Birrer, S., Treu, T., Rusu, C. E., et al. 2019, MNRAS, 484, 4726
- Böker et al. (2022) Böker, T., Arribas, S., Lützgendorf, N., et al. 2022, A&A, 661, A82
- Bolton et al. (2008) Bolton, A. S., Burles, S., Koopmans, L. V. E., et al. 2008, ApJ, 682, 964
- Bolton et al. (2006) Bolton, A. S., Burles, S., Koopmans, L. V. E., Treu, T., & Moustakas, L. A. 2006, ApJ, 638, 703
- Cappellari (2016) Cappellari, M. 2016, ARA&A, 54, 597
- Cappellari (2017) Cappellari, M. 2017, MNRAS, 466, 798
- Cappellari (2023) Cappellari, M. 2023, MNRAS, 526, 3273
- Cappellari & Emsellem (2004) Cappellari, M. & Emsellem, E. 2004, PASP, 116, 138
- Chen et al. (2021) Chen, G. C. F., Fassnacht, C. D., Suyu, S. H., et al. 2021, A&A, 652, A7
- Courteau et al. (2014) Courteau, S., Cappellari, M., de Jong, R. S., et al. 2014, Reviews of Modern Physics, 86, 47
- D’Ago et al. (2023) D’Ago, G., Spiniello, C., Coccato, L., et al. 2023, A&A, 672, A17
- Dressler (1984) Dressler, A. 1984, ApJ, 281, 512
- Dutton et al. (2011) Dutton, A. A., Brewer, B. J., Marshall, P. J., et al. 2011, MNRAS, 417, 1621
- Faber & Jackson (1976) Faber, S. M. & Jackson, R. E. 1976, ApJ, 204, 668
- Falcón-Barroso et al. (2011) Falcón-Barroso, J., Sánchez-Blázquez, P., Vazdekis, A., et al. 2011, A&A, 532, A95
- Jakobsen et al. (2022) Jakobsen, P., Ferruit, P., Alves de Oliveira, C., et al. 2022, A&A, 661, A80
- Kelson et al. (2000) Kelson, D. D., Illingworth, G. D., van Dokkum, P. G., & Franx, M. 2000, ApJ, 531, 159
- Knabel et al. (2024) Knabel, S., Treu, T., Cappellari, M., et al. 2024, arXiv e-prints, arXiv:2409.10631
- Knowles et al. (2021) Knowles, A. T., Sansom, A. E., Allende Prieto, C., & Vazdekis, A. 2021, MNRAS, 504, 2286
- Knowles et al. (2023) Knowles, A. T., Sansom, A. E., Vazdekis, A., & Allende Prieto, C. 2023, MNRAS, 523, 3450
- Koopmans et al. (2003) Koopmans, L. V. E., Treu, T., Fassnacht, C. D., Blandford, R. D., & Surpi, G. 2003, ApJ, 599, 70
- Mamon & Łokas (2005) Mamon, G. A. & Łokas, E. L. 2005, MNRAS, 363, 705
- Mehrgan et al. (2023) Mehrgan, K., Thomas, J., Saglia, R., Parikh, T., & Bender, R. 2023, ApJ, 948, 79
- Morrissey et al. (2018) Morrissey, P., Matuszewski, M., Martin, D. C., et al. 2018, The Astrophysical Journal, 864, 93
- Sánchez-Blázquez et al. (2006) Sánchez-Blázquez, P., Peletier, R. F., Jiménez-Vicente, J., et al. 2006, MNRAS, 371, 703
- Schwarz (1978) Schwarz, G. 1978, The Annals of Statistics, 6
- Shajib et al. (2022) Shajib, A. J., Glazebrook, K., Barone, T., et al. 2022, The Astrophysical Journal, 938, 141
- Shajib et al. (2023) Shajib, A. J., Mozumdar, P., Chen, G. C. F., et al. 2023, A&A, 673, A9
- Shajib et al. (2018) Shajib, A. J., Treu, T., & Agnello, A. 2018, MNRAS, 473, 210
- Shu et al. (2015) Shu, Y., Bolton, A. S., Brownstein, J. R., et al. 2015, ApJ, 803, 71
- Spiniello et al. (2021a) Spiniello, C., Tortora, C., D’Ago, G., et al. 2021a, A&A, 646, A28
- Spiniello et al. (2021b) Spiniello, C., Tortora, C., D’Ago, G., et al. 2021b, A&A, 654, A136
- Spiniello et al. (2015) Spiniello, C., Trager, S. C., & Koopmans, L. V. E. 2015, ApJ, 803, 87
- Suyu et al. (2010) Suyu, S. H., Marshall, P. J., Auger, M. W., et al. 2010, ApJ, 711, 201
- Treu (2010) Treu, T. 2010, ARA&A, 48, 87
- Treu & Koopmans (2002) Treu, T. & Koopmans, L. V. E. 2002, ApJ, 575, 87
- Treu & Marshall (2016) Treu, T. & Marshall, P. J. 2016, A&A Rev., 24, 11
- Treu et al. (2001) Treu, T., Stiavelli, M., Møller, P., Casertano, S., & Bertin, G. 2001, MNRAS, 326, 221
- Treu et al. (2022) Treu, T., Suyu, S. H., & Marshall, P. J. 2022, Strong lensing time-delay cosmography in the 2020s
- Valdes et al. (2004) Valdes, F., Gupta, R., Rose, J. A., Singh, H. P., & Bell, D. J. 2004, The Astrophysical Journal Supplement Series, 152, 251
- Vazdekis et al. (2016) Vazdekis, A., Koleva, M., Ricciardelli, E., Röck, B., & Falcón-Barroso, J. 2016, Monthly Notices of the Royal Astronomical Society, 463, 3409
- Vazdekis et al. (2010) Vazdekis, A., Sánchez-Blázquez, P., Falcón-Barroso, J., et al. 2010, MNRAS, 404, 1639
- Verro et al. (2022) Verro, K., Trager, S. C., Peletier, R. F., et al. 2022, A&A, 660, A34
- Verro et al. (2022) Verro, K., Trager, S. C., Peletier, R. F., et al. 2022, Astronomy and Astrophysics, 661, A50
- Westfall et al. (2019) Westfall, K. B., Cappellari, M., Bershady, M. A., et al. 2019, AJ, 158, 231
- Yıldırım et al. (2020) Yıldırım, A., Suyu, S. H., & Halkola, A. 2020, Monthly Notices of the Royal Astronomical Society, 493, 4783
Appendix A Derivation of weighted combination formulae

### v2

@@ Acknowledgements.
The first two authors should be regarded as joint first authors. We are grateful to S. Trager and A. Vazdekis for sharing with us their quality flags of the XSL and MILES libraries, respectively. We thank A.S. Bolton and M.Bernardi for useful discussions regarding SDSS velocity dispersions. We thank E. Buckley-Geer, M. Millon, and all the friends of the TDCOSMO collaboration for useful feedback that improved this manuscript. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1974. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. Based in part on observations collected at the European Southern Observatory data obtained from the ESO Science Archive Facility. Some of the data presented here were taken from the Sloan Digital Sky Survey Archive. Support for this work was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51492 awarded to AJS by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. AJS also received support from NASA through the STScI grants HST-GO-16773 and JWST-GO-2974. We acknowledge support by NSF through grants NSF-AST-1906976, NSF-AST-2407277, and NSF-AST-1836016, and from the Moore Foundation through grant 8548. We acknowledge support from NASA through grants JWST-GO-1974 and JWST-GO-2974. SB thanks the Department of Physics and Astronomy, Stony Brook, for their support.Appendix A Comparisons of template libraries for various samples
Here, we show the comparison of velocity dispersions across template libraries as in Figure 10 for the KCWI, SDSS, and JWST/NIRSpec samples in Figures 14 to 16.
Appendix B Validation on mock data
We validated the accuracy of the methods described in Section 5 by applying the recipe to a mock dataset. We generated two samples of mock spectra. In the first, mock spectra are based on stellar templates from XSL-DR3 and designed to simulate spectra from a KCWI observation. The resolution of XSL templates (13 km s-1) is better than the target resolution of KCWI data (35 km s-1), so the effect of the instrumental line spread function is simulated. We simulated KCWI data because the sample’s high S/N (mean 160 Å-1) allows us to probe the accuracy of our methods over a range of realistic noise levels. The second test is identical to the first except for the following: the mock spectra are based on templates from the Indo-US library at the native template resolution of km s-1.
To begin, we selected a representative object from our KCWI galaxy sample, SDSSJ1204+0358. This object has a typical velocity dispersion for lens galaxies ( km s-1). From the fiducial fits to this object using both the XSL and Indo-US libraries, in the wavelength range 3600-4500 with additive and multiplicative polynomials of 6 and 2, we collected the weights of the templates used for the fit. The XSL fit utilized 13 of the available 496 templates with normalized weights greater than 1%, while the Indo-US fit utilized 16 of the available 989 templates. These templates were log-sampled in wavelength. For the mock XSL sample, templates are convolved to the resolution of KCWI. From the subset of 13 (16) templates that compose the original combination in the XSL (Indo-US) fit, we constructed three new mock spectra. We randomly alter the weights of each template by 10% and take a spectrum that results in a root-mean-square residual (averaged across wavelengths) ¡ 10% compared with the original combination, to avoid unrealistic alterations to the stellar population represented by the linear combination of templates. We oversampled the spectrum by a factor of 12 to ensure an accurate convolution with the Gaussian LOSVD, which has a mean velocity of zero and a velocity dispersion of 300 km s-1. The resulting convolved spectrum is rebinned back to KCWI resolution. For each of the three mock spectra, we generated 100 noise realizations for each of seven different S/N per : 15, 30, 45, 60, 90, 120, and 150, by dividing the flux of the spectra by the desired S/N, and by the square root of the KCWI pixel size () for the XSL sample. We tested the recipe outlined in Section 5 on these mock samples, each consisting of 2100 spectra. Our goal is to assess the method’s ability to constrain the systematic uncertainties introduced by the template libraries on three bases, determining whether:
1.
Our method can correctly identify the most appropriate source of templates, namely, the parent template library used to generate the mock sample;
2.
The resulting weighted averages result in the correct velocity dispersion within ;
3.
The S/N threshold of 30 per results in systematic errors within 1, as we have found for our observed datasets.
For each mock spectrum, we fit in the wavelength range 3600-4500 with the clean Indo-US, MILES, and XSL libraries. We did not introduce polynomial corrections to the continuum, since they were not included in the mock galactic spectra. Figure 2 shows that there is a mismatch in continuum shape between the template libraries for the same observed stars. For the mock XSL sample, the correction is not needed for fits that utilize XSL templates, and a high polynomial order could bias the resulting velocity dispersion while at the same time not being high enough to properly account for the Indo-US and MILES fits. The resolutions of Indo-US and MILES are lower than the mock XSL spectra, so the templates cannot be broadened to the data resolution. Instead, the correction is applied in quadrature after the measurement utilizing Equation 1. For the mock Indo-US spectra, the same post-measurement correction is applied only for MILES. For each sample and each S/N, we combine the subsample of 300 mock spectra with and without the BIC-weighting methods described in Section 5. The Bessel correction is implemented for both weighted and unweighted systematic uncertainties.
Weights were calculated from the across each S/N subsample and are given for the XSL and Indo-US mock samples in Tables 3 and 4, respectively. For the XSL mock sample, on average across the S/N subsamples, the BIC-weighting prefers XSL with 99.97 of the weight, with the other 0.03 going to Indo-US. This strong preference is expected. On the other hand, the Indo-US mock sample only mildly prefers its parent template library, with almost equal weight given to MILES and about half that weight given to XSL for all subsamples with S/N per . This might be because the mock spectra are sampled at lower resolution than the XSL mock sample.
Results for the XSL and Indo-US mock samples are listed in Tables 5 and 6 shown in Figures 17 and 18, where the left panel shows the weighted results and the right panel shows the unweighted results. In both panels, the x-axis shows the S/N per , and the y-axis shows the fractional deviation from the true value of 300 km s-1. The red, blue, green markers show the unweighted mean values for Indo-US, MILES, and XSL fits, respectively, for the given S/N subsample. Black markers show the mean value over the template libraries combined over the whole subsample, with two error bars each, representing the systematic errors with narrow horizontal cap-width and the statistical errors with the comparatively wider horizontal cap-width.
For both mock samples, the weighted mean velocity dispersions are within of the mean for all S/N subsamples, indicating that the method does consistently return sub-percent accuracy. For the XSL mock sample, the unweighted mean velocity dispersions are unable to improve accuracy to within 1%, regardless of S/N. Fits to the XSL mock sample with Indo-US and MILES libraries overestimate the true velocity dispersion by , which is likely because they might require some polynomial corrections. On the other hand, the Indo-US mock sample is accurate to within 1% for the unweighted case as well.
In terms of estimating the systematic errors, the XSL mock sample results in slightly overestimated systematic uncertainties for the weighted case in comparison with the unweighted case. This is a result of the extreme imbalance in the weights, which inflates the uncertainty through the Bessel correction (see Equation 14), where the term explodes when the sum of weights approaches 1. This is highly unlikely to occur for real data, where our results clearly show sub-percent systematic uncertainties (see Section 6). Even in this odd scenario, the unweighted and weighted estimate of the systematic errors reach 1.2 and 1.25%, and the BIC-weighting is important for ensuring sub-percent accuracy. The Indo-US mock sample results in sub-percent estimates of the systematic uncertainties for all S/N. Across the subsamples, there is a slight improvement in the estimate over the unweighted case.
In summary, our tests with mock spectra indicate the following:
1.
Our method returns the true velocity dispersion to within 1% accuracy for all S/N subsamples when utilizing the BIC weights and approaches this threshold even without using the weights;
2.
Our method correctly identifies the most appropriate set of stellar templates, the parent library from which the mock spectra are generated;
3.
For S/N per and above the results are accurate within 1%, as we find in real data. The weighted estimate of the systematic error is within 1% for the Indo-US mock, while it is slightly above 1% for the XLS mock, owing to numerical issues with the Bessell correction when the weight of the preferred library is so close to unity.
| S/N per | Indo-US | MILES | XSL |
| 15 | 0.0006 | 0.0001 | 0.9994 |
| 30 | 0.0000 | 0.0000 | 1.0000 |
| 45 | 0.0000 | 0.0000 | 1.0000 |
| 60 | 0.0000 | 0.0000 | 1.0000 |
| 90 | 0.0000 | 0.0000 | 1.0000 |
| 120 | 0.0003 | 0.0000 | 0.9997 |
| 150 | 0.0015 | 0.0000 | 0.9985 |
| S/N per | Indo-US | MILES | XSL |
| 15 | 0.6587 | 0.0057 | 0.3355 |
| 30 | 0.4635 | 0.3048 | 0.2317 |
| 45 | 0.4065 | 0.3902 | 0.2033 |
| 60 | 0.4038 | 0.3944 | 0.2019 |
| 90 | 0.4016 | 0.3976 | 0.2008 |
| 120 | 0.4012 | 0.3981 | 0.2006 |
| 150 | 0.4012 | 0.3982 | 0.2006 |
| S/N per | ||||||
| 15 | -0.11% | 1.65% | 3.73% | 1.29% | 1.52% | 3.88% |
| 30 | 0.05% | 1.37% | 1.93% | 1.38% | 1.30% | 2.08% |
| 45 | -0.16% | 1.29% | 1.30% | 1.13% | 1.20% | 1.46% |
| 60 | -0.25% | 1.27% | 0.98% | 1.07% | 1.22% | 1.16% |
| 90 | -0.28% | 1.25% | 0.67% | 1.03% | 1.22% | 0.89% |
| 120 | -0.28% | 1.25% | 0.51% | 1.05% | 1.22% | 0.77% |
| 150 | -0.27% | 1.26% | 0.40% | 1.06% | 1.23% | 0.70% |
| S/N per | ||||||
| 15 | 1.05% | 0.88% | 4.01% | 0.77% | 1.01% | 4.06% |
| 30 | 0.80% | 0.58% | 2.15% | 0.78% | 0.59% | 2.14% |
| 45 | 0.47% | 0.45% | 1.50% | 0.51% | 0.47% | 1.48% |
| 60 | 0.47% | 0.37% | 1.18% | 0.52% | 0.39% | 1.16% |
| 90 | 0.35% | 0.38% | 0.86% | 0.43% | 0.41% | 0.84% |
| 120 | 0.31% | 0.39% | 0.72% | 0.40% | 0.43% | 0.69% |
| 150 | 0.35% | 0.40% | 0.65% | 0.44% | 0.44% | 0.61% |
Appendix C Derivation of weighted combination formulae